# Characterizing the Distinguishability of Product Distributions through Multicalibration

### v3

@@ 5 Defining One Intermediate Variable
    Random variable $\widehat{X}_{0}$: Consider the multicalibrated partition $\mathcal{P}=\{P_{i}\}$ guaranteed by the Multicalibration Theorem when applied to the function $g$, distribution $\mathcal{D}$ over domain $\mathcal{X}$, class of functions $\mathcal{F}^{\prime}$, and parameters $\varepsilon^{\prime}$ and $\gamma=(\varepsilon^{\prime})^{2}$ to obtain a $(\mathcal{F}^{\prime},\varepsilon^{\prime},\gamma)$-approximately multicalibrated partition $\mathcal{P}$ of $\mathcal{X}$. Given the multicalibrated partition, construct the random variable $\widehat{X}_{0}$ as follows: to sample according to $\widehat{X}_{0}$, first choose a piece of the partition $P_{i}\in\mathcal{P}$, where $P_{i}$ is chosen with probability $\mathbb{P}[X_{0}\in P_{i}]$. If $\alpha_{1}(P_{i})\geq\sqrt{\varepsilon}$, then sample $x\in P_{i}$ according to $X_{1}|_{P_{i}}$. Otherwise, sample $x$ according to $X_{0}|_{P_{i}}$. This final conditioning step is the key difference from Definition 1.7.
@@
    $X_{0}^{\otimes k}$ and $X_{1}^{\otimes k}$ are $d_{\mathrm{TV}}((\widehat{X}_{0})^{\otimes k},X_{1}^{\otimes k})-2k\varepsilon$-distinguishable by functions in
@@
    with $K_{1}=O(k\cdot\log(|\mathcal{X}|/\varepsilon)\cdot\log(|\mathcal{X}|)/\varepsilon^{12})+O(k\log(|\mathcal{X}|)\cdot\mathrm{polylog}(1/\varepsilon))$, and $K_{2}=O(k/\varepsilon^{4})$.
@@ 5.1 Computational Indistinguishability
Let us focus on the first on the summation over $P_{i}\in\mathcal{P}_{1}$. Observe that on all $P_{i}\in\mathcal{P}_{1}$, by definition $\widetilde{X}^{(\mathcal{U})}_{0}|_{P_{i}}=\mathcal{U}|_{P_{i}}$. This gives the following:
@@
$=\sum_{P_{i}\in\mathcal{P}_{1}}\mathbb{P}[X_{0}\in P_{i}]\left|\sum_{x\in P_{i}}f(x)\cdot\left(\mathbb{P}[X_{0}|_{P_{i}}=x]-\mathbb{P}[U|_{P_{i}}=x]\right)\right|$
$=\sum_{P_{i}\in\mathcal{P}_{1}}\alpha_{0}(P_{i})\cdot 2\mathbb{P}[\mathcal{D}\in P_{i}]\left|\sum_{x\in P_{i}}f(x)\cdot\left(\mathbb{P}[X_{0}|_{P_{i}}=x]-\mathbb{P}[U|_{P_{i}}=x]\right)\right|$
Using Lemma 3.3 to bound the indistinguishability of $X_{0}|_{P_{i}}$ and $U|_{P_{i}}$, we obtain:
@@
[Theorem 5.2] For $X_{0},X_{1},\mathcal{F},\varepsilon>0$, define $\widehat{X}_{0}$ as in Definition 5.1. For $f\in\mathcal{F}^{\otimes k}$, the following three inequalities hold:
@@
$+\left|\sum_{\bar{x}\in\mathcal{X}^{k}}f(\bar{x})\cdot\left(\mathbb{P}\left[(\widehat{X}_{0})^{\otimes k}=\bar{x}\right]-\mathbb{P}\left[X_{1}^{\otimes k}=\bar{x}\right]\right)\right|$
@@ 5.2 Moving from Total Variation Distance to Partitions
$=\sum_{\begin{subarray}{c}P_{i}\in\mathcal{P}\\ \alpha_{1}(P_{i})\geq\varepsilon\end{subarray}}\sum_{\begin{subarray}{c}x\in P_{i}\\ X_{1}(x)>\widehat{X}_{0}(x)\end{subarray}}\left(X_{1}(x)-\widehat{X}_{0}(x)\right)+\sum_{\begin{subarray}{c}P_{i}\in\mathcal{P}\\ \alpha_{1}(P_{i})<\varepsilon\end{subarray}}\sum_{\begin{subarray}{c}x\in P_{i}\\ X_{1}(x)>\widehat{X}_{0}(x)\end{subarray}}\left(X_{1}(x)-\widehat{X}_{0}(x)\right).$
@@
$=\sum_{\begin{subarray}{c}P^{(1)},\dots P^{(k)}\in\mathcal{P}^{\otimes k}\\ \alpha_{1}(P^{(i)})\geq\varepsilon\end{subarray}}\sum_{\begin{subarray}{c}(x_{1},\dots x_{k}):x_{i}\in P^{(i)}\\ X_{1}^{\otimes k}(x_{1},\dots x_{k})>\left(\widehat{X}_{0}\right)^{\otimes k}(x_{1},\dots x_{k})\end{subarray}}\left(X_{1}^{\otimes k}(x_{1},\dots x_{k})-\left(\widehat{X}_{0}\right)^{\otimes k}(x_{1},\dots x_{k})\right)$
$+\sum_{\begin{subarray}{c}P^{(1)},\dots P^{(k)}\in\mathcal{P}^{\otimes k}\\ \exists i:\alpha_{1}(P^{(i)})<\varepsilon\end{subarray}}\sum_{\begin{subarray}{c}(x_{1},\dots x_{k}):x_{i}\in P^{(i)}\\ X_{1}^{\otimes k}(x_{1},\dots x_{k})>\left(\widehat{X}_{0}\right)^{\otimes k}(x_{1},\dots x_{k})\end{subarray}}\left(X_{1}^{\otimes k}(x_{1},\dots x_{k})-\left(\widehat{X}_{0}\right)^{\otimes k}(x_{1},\dots x_{k})\right).$
@@ 5.3 Efficient Distinguishers
with $K_{1}=O(k/\varepsilon^{12}\cdot\log(|\mathcal{X}|)\cdot\log(|\mathcal{X}|/\varepsilon))+O(k\mathrm{polylog}(1/\varepsilon\delta))$, and $K_{2}=O(k/\varepsilon^{4})$, which distinguishes $X^{\otimes k}_{0},X_{1}^{\otimes k}$ with advantage $d_{\mathrm{TV}}((\widehat{X}_{0})^{\otimes k},X_{1}^{\otimes k})-4\delta k-2k\varepsilon$.
@@
where $q=O(1/\varepsilon^{\prime 2}),r=O(1/\varepsilon^{\prime}),t=O(1/(\varepsilon^{\prime 4}\gamma)\cdot\log(|\mathcal{X}|)/\varepsilon^{\prime})$. By plugging in $\varepsilon^{\prime}=\varepsilon^{2}$, we see that this circuit lies in
@@
where $K_{1}=O(k/\varepsilon^{12}\cdot\log(|\mathcal{X}|)\cdot\log(|\mathcal{X}|/\varepsilon))+O(k\mathrm{polylog}(1/\varepsilon\delta))$, and $K_{2}=O(k/\varepsilon^{4})$.
@@
where we used 5.7 in the final step.
@@ 6 Re-deriving a Result of Halevi and Rabin
′′$\mathcal{F}^{\prime\prime}:=\mathcal{F}^{\prime}_{K_{1},K_{2}},$
with $K_{1}=O(\frac{1}{\varepsilon^{6}}\cdot\log(|\mathcal{X}|)/\varepsilon)+O(\mathrm{polylog}(1/\varepsilon))$, and $K_{2}=O(1/\varepsilon^{2})$.
@@ 7 Characterizing Distinguishability through Pseudo-Hellinger and Pseudo-Rényi Distance
    $X_{0}^{\otimes k},X_{1}^{\otimes k}$ are $(1-e^{-k\delta^{2}}-2k\varepsilon)$ distinguishable by functions $f\in\mathcal{F}_{K_{2},K_{1}}$, where $K_{1}=O(\frac{k}{\varepsilon^{6}}\cdot\log(|\mathcal{X}|)\cdot\log(|\mathcal{X}|/\varepsilon))+O(k\cdot\log(|\mathcal{X}|)\cdot\mathrm{polylog}(k/\varepsilon))$ and $K_{2}=O(k/\varepsilon^{2})$.
@@
    $X_{0}^{\otimes k},\mathcal{U}^{\otimes k}$ are $(1-e^{-\Omega(k\min\{\delta,1\})}-2k\varepsilon)$ distinguishable by functions $f\in\mathcal{F^{\prime}}_{K_{2},K_{1}}$ with $K_{1}=O(k\cdot\log(|\mathcal{X}|/\varepsilon)\cdot\log(|\mathcal{X}|)/\varepsilon^{12})+O(k\log(|\mathcal{X}|)\cdot\mathrm{polylog}(1/\varepsilon))$, and $K_{2}=O(k/\varepsilon^{4})$.
@@
To prove Part (1), following the same steps as in the proof of Theorem 7.1, by Definition 1.2 and a hybrid argument we can bound the computational indistinguishability of $X_{0}^{\otimes k}$ and $\mathcal{U}^{\otimes k}$ by the computational indistinguishability of $\widetilde{X}_{0}^{\otimes k}$ and $\mathcal{U}^{\otimes k}$, plus $2k\varepsilon$. This is upper-bounded by the total variation distance between $\widetilde{X}_{0}^{\otimes k}$ and $\mathcal{U}^{\otimes k}$, plus $2k\varepsilon$, which is itself bounded above by $\sqrt{2kd_{H}^{2}\left(\widetilde{X}_{0},\mathcal{U}\right)}+2k\varepsilon$. The bound in Part (1) of the theorem is then obtained by applying 2.2, which tells us $d_{H}^{2}\left(\widetilde{X}_{0},\mathcal{U}\right)=1-\sqrt{2^{-\delta}}\leq\sqrt{\delta}$.
To prove Part (2), again as in the proof of Theorem 7.1, use Theorem 1.1 (more specifically, its generalization in Theorem 5.1) to obtain a $\widetilde{X}_{0}$ that is $\varepsilon$-indistinguishable from $X_{0}$ such that the computational distinguishability of $X_{0}^{\otimes k},\mathcal{U}^{\otimes k}$ by functions $f\in\mathcal{F}_{K_{2},K_{1}}$ is at least $d_{\mathrm{TV}}\left(\widetilde{X}_{0}^{\otimes k},\mathcal{U}^{\otimes k}\right)-2k\varepsilon$. As before, this is at least $1-e^{-k\cdot d^{2}_{\mathrm{H}}(\widetilde{X}_{0},\mathcal{U})}-2k\varepsilon$, which can be related to the Rényi entropy of order $1/2$ of $\widetilde{X}_{0}$ using 2.2. 2.2 tells us $d_{H}^{2}\left(\widetilde{X}_{0},\mathcal{U}\right)=1-\sqrt{2^{-\delta}}$, which is $\Omega(\min\{1,\delta\})$, giving the theorem. ∎
@@ 8 General Product Distributions
    $Y^{1}_{0}\otimes\dots\otimes Y^{k}_{0}$ and $Y^{1}_{1}\otimes\dots\otimes Y^{k}_{1}$ are $(d_{\mathrm{TV}}(\widetilde{Y}^{1}_{0}\otimes\dots\otimes\widetilde{Y}^{k}_{0},\widetilde{Y}^{1}_{1}\otimes\dots\otimes\widetilde{Y}^{k}_{1})-2k\cdot\varepsilon)$-distinguishable by functions in $\mathcal{F^{\prime}}_{K_{2},K_{1}}$, where $K_{1}=O(\frac{k}{\varepsilon^{6}}\cdot\log(|\mathcal{X}|)\cdot\log(|\mathcal{X}|/\varepsilon))+O(k\cdot\log(|\mathcal{X}|)\cdot\mathrm{polylog}(k/\varepsilon))$ and $K_{2}=O(k/\varepsilon^{2})$.
@@
$+\left|\sum_{\bar{x}\in\mathcal{X}^{k}}f(\bar{x})\cdot\left(\mathbb{P}[\widetilde{Y}^{1}_{0}\otimes\dots\otimes\widetilde{Y}^{k}_{0}=\bar{x}]-\mathbb{P}[\widetilde{Y}^{1}_{1}\otimes\dots\otimes\widetilde{Y}^{k}_{1}=\bar{x}]\right)\right|$
$+\left|\sum_{\bar{x}\in\mathcal{X}^{k}}f(\bar{x})\cdot\left(\mathbb{P}[\widetilde{Y}^{1}_{1}\otimes\dots\otimes\widetilde{Y}^{k}_{1}=\bar{x}]-\mathbb{P}[Y^{1}_{1}\otimes\dots\otimes Y^{k}_{1}=\bar{x}]\right)\right|.$
@@
where $K_{1}=O(\frac{k}{\varepsilon^{6}}\cdot\log(|\mathcal{X}|)\cdot\log(|\mathcal{X}|/\varepsilon))+O(k\cdot\log(|\mathcal{X}|)\cdot\mathrm{polylog}(k/\varepsilon))$ and $K_{2}=O(k/\varepsilon^{2})$, which distinguishes $Y^{1}_{0}\otimes\dots\otimes Y_{0}^{k},Y^{1}_{1}\otimes\dots\otimes Y_{1}^{k}$ with advantage
@@
For a sequence of $k$ samples $x_{1},\dots x_{k}$, the optimal distinguisher is simply the distinguisher which returns $1$ (corresponding to deciding that the samples are from $\widetilde{X}_{1,\delta}^{\otimes k}$) if the above expression is $\geq 1$, and 00 otherwise (corresponding to deciding that the samples are from $\widetilde{X}_{0,\delta}^{\otimes k}$). This is because in general for two distributions $p,q$, we have that
@@
$=d_{\mathrm{TV}}(\widetilde{Y}^{1}_{0,\delta}\otimes\dots\otimes\widetilde{Y}_{0,\delta}^{k},\widetilde{Y}^{1}_{1,\delta}\otimes\dots\otimes\widetilde{Y}_{1,\delta}^{k}).$
Now, all that remains to show is that we can compute such a circuit $C$ which is not too complex. Indeed, the key observation is that because we “rounded” our distributions initially, now when we compute the product $\prod_{a=1}^{k}\frac{\alpha^{(a)}_{1,\delta}(P_{i_{j}})}{\alpha^{(a)}_{0,\delta}(P_{i_{j}})}$, every pair of $\alpha^{(a)}_{1,\delta}(P_{i_{j}}),\alpha^{(a)}_{0,\delta}(P_{i_{j}})$ share a denominator. Thus, computing whether this product is $\geq 1$ is equivalent to multiplying together $k$ integers in $[r/\delta]$ and checking which is larger. So, we consider the following circuit: first, $C$ has $k$ copies of the multicalibration circuit which, upon receiving element $x_{a}$, computes the index $i$ for which $x_{a}\in P^{(a)}_{i}$ (defined with respect to the $a$th distribution). Recall that this circuit is $\in\mathcal{F}^{\prime}_{t,q,r}$. Next, once we have computed the index $i$ of the partition that an element $x$ is in, we hard-wire two-lookup tables: one computes the numerator of $\alpha^{(a)}_{0,\delta}(P_{i})$, and the other computes the numerator of $\alpha^{(a)}_{1,\delta}(P_{i})$. Observe that because there are only $r$ partitions, the complexity of these look-up tables is not too large. By [Lup70], each look-up table requires only circuits of size $O(r)$. Once we have computed the numerators of these values, all that remains is to multiply together these integers, and return the index $b\in\{0,1\}$, corresponding to whichever product is larger. Computing this product can be done with circuits of size $O(k\cdot\mathrm{polylog}(r/\delta))$.
@@
as we use $k$ copies of a circuit in $\mathcal{F}_{t,q\log(|\mathcal{X}|)}$, along with $O(k)$ look-up tables, and $O(k\cdot\mathrm{polylog}(r/\delta))$ extra wires to compute the product. By plugging in $q=O(1/\varepsilon^{2}),r=O(1/\varepsilon),t=O(1/(\varepsilon^{4}\gamma)\cdot\log(|\mathcal{X}|)/\varepsilon)$, we obtain the stated complexity.

### v5

@@ 5 Defining One Intermediate Variable
    Random variable $\widehat{X}_{0}$: Consider the multicalibrated partition $\mathcal{P}=\{P_{i}\}$ guaranteed by the Multicalibration Theorem when applied to the function $g$, distribution $\mathcal{D}$ over domain $\mathcal{X}$, class of functions $\mathcal{F}^{\prime}$, and parameters $\varepsilon^{\prime}$ and $\gamma=(\varepsilon^{\prime})^{2}$ to obtain a $(\mathcal{F}^{\prime},\varepsilon^{\prime},\gamma)$-approximately multicalibrated partition $\mathcal{P}$ of $\mathcal{X}$. Given the multicalibrated partition, construct the random variable $\widehat{X}_{0}$ as follows: to sample according to $\widehat{X}_{0}$, first choose a piece of the partition $P_{i}\in\mathcal{P}$, where $P_{i}$ is chosen with probability $\mathbb{P}[X_{0}\in P_{i}]$. If $\alpha_{1}(P_{i})\geq\sqrt{\varepsilon^{\prime}}$, then sample $x\in P_{i}$ according to $X_{1}|_{P_{i}}$. Otherwise, sample $x$ according to $X_{0}|_{P_{i}}$. This final conditioning step is the key difference from Definition 1.15.
@@
    $X_{0}^{\otimes k}$ and $X_{1}^{\otimes k}$ are $(d_{\mathrm{TV}}((\widehat{X}_{0})^{\otimes k},X_{1}^{\otimes k})-3k\varepsilon)$-distinguishable by functions in
@@
    with $K_{1}=O(k\cdot\log(|\mathcal{X}|/\varepsilon)\cdot\log(|\mathcal{X}|)/\varepsilon^{24})+O(k\log(|\mathcal{X}|)\cdot\mathrm{polylog}(1/\varepsilon))$, and $K_{2}=O(k/\varepsilon^{12})$.
@@ 5.1 Computational Indistinguishability
Let us focus on the first on the summation over $P_{i}\in\mathcal{P}_{1}$. Observe that on all $P_{i}\in\mathcal{P}_{1}$, by definition $\widehat{X}^{(\mathcal{U})}_{0}|_{P_{i}}=\mathcal{U}|_{P_{i}}$, the uniform distribution over elements of $P_{i}$. This gives the following:
@@
$\sum_{P_{i}\in\mathcal{P}_{1}}\mathbb{P}[X_{0}\in P_{i}]\left|\sum_{x\in P_{i}}f(x)\cdot\left(\mathbb{P}[X_{0}|_{P_{i}}=x]-\mathbb{P}[\mathcal{U}|_{P_{i}}=x]\right)\right|$
$\sum_{P_{i}\in\mathcal{P}_{1}}\alpha_{0}(P_{i})\cdot 2\mathbb{P}[\mathcal{D}\in P_{i}]\left|\sum_{x\in P_{i}}f(x)\cdot\left(\mathbb{P}[X_{0}|_{P_{i}}=x]-\mathbb{P}[\mathcal{U}|_{P_{i}}=x]\right)\right|$
Using Lemma 3.4 to bound the indistinguishability of $X_{0}|_{P_{i}}$ and $\mathcal{U}|_{P_{i}}$, we obtain:
@@
[Lemma 5.3] For $X_{0},X_{1},\mathcal{F},\varepsilon>0$, define $\widehat{X}_{0}$ as in Definition 5.1. For $f\in\mathcal{F}^{\otimes k}$, the following three inequalities hold:
@@
$\left|\sum_{\bar{x}\in\mathcal{X}^{k}}f(\bar{x})\cdot\left(\mathbb{P}\left[(\widehat{X}_{0})^{\otimes k}=\bar{x}\right]-\mathbb{P}\left[X_{1}^{\otimes k}=\bar{x}\right]\right)\right|$
@@ 5.2 Moving from Total Variation Distance to Partitions
$\sum_{\begin{subarray}{c}P_{i}\in\mathcal{P}\\ \alpha_{1}(P_{i})\geq\varepsilon\end{subarray}}\sum_{\begin{subarray}{c}x\in P_{i}\\ X_{1}(x)>\widehat{X}_{0}(x)\end{subarray}}\left(X_{1}(x)-\widehat{X}_{0}(x)\right)+\sum_{\begin{subarray}{c}P_{i}\in\mathcal{P}\\ \alpha_{1}(P_{i})<\varepsilon\end{subarray}}\sum_{\begin{subarray}{c}x\in P_{i}\\ X_{1}(x)>\widehat{X}_{0}(x)\end{subarray}}\left(X_{1}(x)-\widehat{X}_{0}(x)\right).$
@@
$\sum_{\begin{subarray}{c}P^{(1)},\dots P^{(k)}\in\mathcal{P}^{\otimes k}\\ \alpha_{1}(P^{(i)})\geq\varepsilon\end{subarray}}\sum_{\begin{subarray}{c}(x_{1},\dots x_{k}):x_{i}\in P^{(i)}\\ X_{1}^{\otimes k}(x_{1},\dots x_{k})>\left(\widehat{X}_{0}\right)^{\otimes k}(x_{1},\dots x_{k})\end{subarray}}\left(X_{1}^{\otimes k}(x_{1},\dots x_{k})-\left(\widehat{X}_{0}\right)^{\otimes k}(x_{1},\dots x_{k})\right)$
$\sum_{\begin{subarray}{c}P^{(1)},\dots P^{(k)}\in\mathcal{P}^{\otimes k}\\ \exists i:\alpha_{1}(P^{(i)})<\varepsilon\end{subarray}}\sum_{\begin{subarray}{c}(x_{1},\dots x_{k}):x_{i}\in P^{(i)}\\ X_{1}^{\otimes k}(x_{1},\dots x_{k})>\left(\widehat{X}_{0}\right)^{\otimes k}(x_{1},\dots x_{k})\end{subarray}}\left(X_{1}^{\otimes k}(x_{1},\dots x_{k})-\left(\widehat{X}_{0}\right)^{\otimes k}(x_{1},\dots x_{k})\right).$
@@ 5.3 Efficient Distinguishers
with $K_{1}=O(k/\varepsilon^{24}\cdot\log(|\mathcal{X}|)\cdot\log(|\mathcal{X}|/\varepsilon))+O(k\mathrm{polylog}(1/\varepsilon\delta))$, and $K_{2}=O(k/\varepsilon^{12})$, which distinguishes $X^{\otimes k}_{0},X_{1}^{\otimes k}$ with advantage $d_{\mathrm{TV}}((\widehat{X}_{0})^{\otimes k},X_{1}^{\otimes k})-4\delta k-2k\varepsilon$.
@@
where $q=O(1/\varepsilon^{\prime 6}),r=O(1/\varepsilon^{\prime}),t=O(1/(\varepsilon^{\prime 4}\gamma)\cdot\log(|\mathcal{X}|)/\varepsilon^{\prime})$. By plugging in $\varepsilon^{\prime}=\varepsilon^{2}$, we see that this circuit lies in
@@
where $K_{1}=O(k/\varepsilon^{24}\cdot\log(|\mathcal{X}|)\cdot\log(|\mathcal{X}|/\varepsilon))+O(k\mathrm{polylog}(1/\varepsilon\delta))$, and $K_{2}=O(k/\varepsilon^{12})$.
@@
where we used 5.8 in the final step.
@@ 6 Re-deriving a Result of Halevi and Rabin
$\mathcal{F}^{\prime\prime}:=\mathcal{F}^{\prime}_{K_{1},K_{2}},$
with $K_{1}=O(\frac{1}{\varepsilon^{12}}\cdot\log(|\mathcal{X}|)/\varepsilon)+O(\mathrm{polylog}(1/\varepsilon))$, and $K_{2}=O(1/\varepsilon^{6})$.
@@ 7 Characterizing Distinguishability through Pseudo-Hellinger and Pseudo-Rényi Distance
    $X_{0}^{\otimes k},X_{1}^{\otimes k}$ are $(1-e^{-k\delta^{2}}-2k\varepsilon)$ distinguishable by functions $f\in\mathcal{F}_{K_{2},K_{1}}$, where $K_{1}=O(\frac{k}{\varepsilon^{12}}\cdot\log(|\mathcal{X}|)\cdot\log(|\mathcal{X}|/\varepsilon))+O(k\cdot\log(|\mathcal{X}|)\cdot\mathrm{polylog}(k/\varepsilon))$ and $K_{2}=O(k/\varepsilon^{6})$.
@@
    $X_{0}^{\otimes k},\mathcal{U}^{\otimes k}$ are $(1-e^{-\Omega(k\min\{\delta,1\})}-2k\varepsilon)$ distinguishable by functions $f\in\mathcal{F^{\prime}}_{K_{2},K_{1}}$ with $K_{1}=O(k\cdot\log(|\mathcal{X}|/\varepsilon)\cdot\log(|\mathcal{X}|)/\varepsilon^{24})+O(k\log(|\mathcal{X}|)\cdot\mathrm{polylog}(1/\varepsilon))$, and $K_{2}=O(k/\varepsilon^{12})$.
@@
To prove Part (1), following the same steps as in the proof of Theorem 7.2, by Definition 1.6 and a hybrid argument we can bound the computational indistinguishability of $X_{0}^{\otimes k}$ and $\mathcal{U}^{\otimes k}$ by the computational indistinguishability of $\widetilde{X}_{0}^{\otimes k}$ and $\mathcal{U}^{\otimes k}$, plus $2k\varepsilon$. This is upper-bounded by the total variation distance between $\widetilde{X}_{0}^{\otimes k}$ and $\mathcal{U}^{\otimes k}$, plus $2k\varepsilon$, which is itself bounded above by $\sqrt{2kd_{H}^{2}\left(\widetilde{X}_{0},\mathcal{U}\right)}+2k\varepsilon$. The bound in Part (1) of the theorem is then obtained by applying 2.3, which tells us $d_{H}^{2}\left(\widetilde{X}_{0},\mathcal{U}\right)=1-\sqrt{2^{-\delta}}\leq\sqrt{\delta}$.
To prove Part (2), again as in the proof of Theorem 7.2, use Theorem 1.1 (more specifically, its generalization in Theorem 5.2) to obtain a $\widetilde{X}_{0}$ that is $\varepsilon$-indistinguishable from $X_{0}$ such that the computational distinguishability of $X_{0}^{\otimes k},\mathcal{U}^{\otimes k}$ by functions $f\in\mathcal{F}_{K_{2},K_{1}}$ is at least $d_{\mathrm{TV}}\left(\widetilde{X}_{0}^{\otimes k},\mathcal{U}^{\otimes k}\right)-2k\varepsilon$. As before, this is at least $1-e^{-k\cdot d^{2}_{\mathrm{H}}(\widetilde{X}_{0},\mathcal{U})}-2k\varepsilon$, which can be related to the Rényi entropy of order $1/2$ of $\widetilde{X}_{0}$ using 2.3. 2.3 tells us $d_{H}^{2}\left(\widetilde{X}_{0},\mathcal{U}\right)=1-\sqrt{2^{-\delta}}$, which is $\Omega(\min\{1,\delta\})$, giving the theorem. ∎
@@ 8 General Product Distributions
    $Y^{1}_{0}\otimes\dots\otimes Y^{k}_{0}$ and $Y^{1}_{1}\otimes\dots\otimes Y^{k}_{1}$ are $(d_{\mathrm{TV}}(\widetilde{Y}^{1}_{0}\otimes\dots\otimes\widetilde{Y}^{k}_{0},\widetilde{Y}^{1}_{1}\otimes\dots\otimes\widetilde{Y}^{k}_{1})-2k\cdot\varepsilon)$-distinguishable by functions in $\mathcal{F^{\prime}}_{K_{2},K_{1}}$, where $K_{1}=O(\frac{k}{\varepsilon^{12}}\cdot\log(|\mathcal{X}|)\cdot\log(|\mathcal{X}|/\varepsilon))+O(k\cdot\log(|\mathcal{X}|)\cdot\mathrm{polylog}(k/\varepsilon))$ and $K_{2}=O(k/\varepsilon^{6})$.
@@
$\left|\sum_{\bar{x}\in\mathcal{X}^{k}}f(\bar{x})\cdot\left(\mathbb{P}[\widetilde{Y}^{1}_{0}\otimes\dots\otimes\widetilde{Y}^{k}_{0}=\bar{x}]-\mathbb{P}[\widetilde{Y}^{1}_{1}\otimes\dots\otimes\widetilde{Y}^{k}_{1}=\bar{x}]\right)\right|$
$\left|\sum_{\bar{x}\in\mathcal{X}^{k}}f(\bar{x})\cdot\left(\mathbb{P}[\widetilde{Y}^{1}_{1}\otimes\dots\otimes\widetilde{Y}^{k}_{1}=\bar{x}]-\mathbb{P}[Y^{1}_{1}\otimes\dots\otimes Y^{k}_{1}=\bar{x}]\right)\right|.$
@@
where $K_{1}=O(\frac{k}{\varepsilon^{12}}\cdot\log(|\mathcal{X}|)\cdot\log(|\mathcal{X}|/\varepsilon))+O(k\cdot\log(|\mathcal{X}|)\cdot\mathrm{polylog}(k/\varepsilon))$ and $K_{2}=O(k/\varepsilon^{6})$, which distinguishes $Y^{1}_{0}\otimes\dots\otimes Y_{0}^{k},Y^{1}_{1}\otimes\dots\otimes Y_{1}^{k}$ with advantage
@@
For a sequence of $k$ samples $x_{1},\dots x_{k}$, the optimal distinguisher is simply the distinguisher which returns $1$ (corresponding to deciding that the samples are from $\widetilde{X}_{1,\delta}^{\otimes k}$) if the above expression is $\geq 1$, and 0 otherwise (corresponding to deciding that the samples are from $\widetilde{X}_{0,\delta}^{\otimes k}$). This is because in general for two distributions $p,q$, we have that
@@
$d_{\mathrm{TV}}(\widetilde{Y}^{1}_{0,\delta}\otimes\dots\otimes\widetilde{Y}_{0,\delta}^{k},\widetilde{Y}^{1}_{1,\delta}\otimes\dots\otimes\widetilde{Y}_{1,\delta}^{k}).$
Now, all that remains to show is that we can compute such a circuit $C$ which is not too complex. Indeed, the key observation is that because we “rounded” our distributions initially, now when we compute the product $\prod_{a=1}^{k}\frac{\alpha^{(a)}_{1,\delta}(P_{i_{j}})}{\alpha^{(a)}_{0,\delta}(P_{i_{j}})}$, every pair of $\alpha^{(a)}_{1,\delta}(P_{i_{j}}),\alpha^{(a)}_{0,\delta}(P_{i_{j}})$ share a denominator. Thus, computing whether this product is $\geq 1$ is equivalent to multiplying together $k$ integers in $r/\delta]$ and checking which is larger. So, we consider the following circuit: first, $C$ has $k$ copies of the multicalibration circuit which, upon receiving element $x_{a}$, computes the index $i$ for which $x_{a}\in P^{(a)}_{i}$ (defined with respect to the $a$th distribution). Recall that this circuit is $\in\mathcal{F}^{\prime}_{t,q,r}$. Next, once we have computed the index $i$ of the partition that an element $x$ is in, we hard-wire two-lookup tables: one computes the numerator of $\alpha^{(a)}_{0,\delta}(P_{i})$, and the other computes the numerator of $\alpha^{(a)}_{1,\delta}(P_{i})$. Observe that because there are only $r$ partitions, the complexity of these look-up tables is not too large. By [Lup70], each look-up table requires only circuits of size $O(r)$. Once we have computed the numerators of these values, all that remains is to multiply together these integers, and return the index $b\in\{0,1\}$, corresponding to whichever product is larger. Computing this product can be done with circuits of size $O(k\cdot\mathrm{polylog}(r/\delta))$.
@@
as we use $k$ copies of a circuit in $\mathcal{F}_{t,q\log(|\mathcal{X}|)}$, along with $O(k)$ look-up tables, and $O(k\cdot\mathrm{polylog}(r/\delta))$ extra wires to compute the product. By plugging in $q=O(1/\varepsilon^{6}),r=O(1/\varepsilon),t=O(1/(\varepsilon^{12}\gamma)\cdot\log(|\mathcal{X}|)/\varepsilon)$, we obtain the stated complexity.